OpenStudy (anonymous):

A 45- 45- 90 triangle has a hypotenuse of length 14 units. What is the length of one of the legs? If necessary, round your answer to two decimal places.

OpenStudy (jdoe0001):

45-45-90 rule http://nanharbison.com/sats/math_images/45-45-90.gif

OpenStudy (anonymous):

Can you draw this triangle?

OpenStudy (ivettef365):

leg 1 = x leg 2 = x hyp = x\(\sqrt{2}\)

OpenStudy (anonymous):

didnt give me a triangle just asled me the question

OpenStudy (anonymous):

@ivettef365 can i get the answer in units

OpenStudy (anonymous):



OpenStudy (anonymous):

9.90?

OpenStudy (anonymous):

and:  A = \[\frac{ \sqrt{2} }{ 2 } X\] and B = \[\frac{ \sqrt{2} }{ 2 } X\] so A+B ?

OpenStudy (anonymous):

got 9.90 was correct

OpenStudy (anonymous):

yes I think

OpenStudy (ivettef365):

just divide 14 /\(\sqrt{2}\) what do you get ?

OpenStudy (anonymous):

@ivettef365 it was 9.90 thank you guys

OpenStudy (ivettef365):

correct, that is what I got